# StructLM: Towards Building Generalist Models for Structured Knowledge Grounding

### 26 Feb 2024 v1

@@ Abstract
Structured data sources, such as tables, graphs, and databases, are ubiquitous knowledge sources. Despite the demonstrated capabilities of large language models (LLMs) on plain text, their proficiency in interpreting and utilizing structured data remains limited. Our investigation reveals a notable deficiency in LLMs’ ability to process structured data, e.g., ChatGPT lags behind state-of-the-art (SoTA) model by an average of 35%. To augment the Structured Knowledge Grounding (SKG) capabilities in LLMs, we have developed a comprehensive instruction tuning dataset comprising 1.1 million examples. Utilizing this dataset, we train a series of models, referred to as StructLM, based on the Code-LLaMA architecture, ranging from 7B to 34B parameters. Our StructLM series surpasses task-specific models on 14 out of 18 evaluated datasets and establishes new SoTA achievements on 7 SKG tasks. Furthermore, StructLM demonstrates exceptional generalization across 6 novel SKG tasks. Contrary to expectations, we observe that scaling model size offers marginal benefits, with StructLM-34B showing only slight improvements over StructLM-7B. This suggests that structured knowledge grounding is still a challenging task and requires more innovative design to push to a new level.
@@ 1 Introduction
The previous work Yu et al. (2020); Liu et al. (2021); Xie et al. (2022); Zhang et al. (2023) has been mostly focused on building task-specific models for different tasks with rather limited generalization ability. Building a generalist structure knowledge grounding (SKG) system across a wide range of tasks proves to be challenging. This is mainly due to the heterogeneity of data format and use cases. We evaluated GPT-3.5-Turbo Jiang et al. (2023) on 18 SKG tasks and observed that its performance is on average 35% lower than the SoTA specialized models. It shows that the LLM’s ability on SKG is heavily overlooked during the pre-training phase.
@@
In this paper, we explore the possibility of building a generalist model based on LLMs that can ground on diverse types of structure and unstructured knowledge to interface with humans. Specifically, we construct a large data set of over a million instruction-following examples, a majority of which is SKG data, along with additional general instruction-following data, which we find improves generalizability. We fine-tune models at three scales: 7B, 13B, and 34B, based on the Code-LLaMA family of code foundation models. When compared to USKG, we find that our 7B model surpasses these single-task models from on $11$ of $18$ tasks with our 34B model outperforming on $14$ of $18$. As depicted in Figure 1, StructLM achieves SoTA on $7$ out of 18 evaluated tasks, beating ChatGPT by a huge margin.
@@ 2.1 Solving SKG tasks
More recently, methods have begun to move away from these auxiliary task-specific structures. USKG Xie et al. (2022) were the first to unify many SKG tasks into a sequence-to-sequence format, allowing them to be aggregated into the same data mixture. StructGPT Jiang et al. (2023) represents a line of work that uses prompting frameworks on powerful LLMs to solve tasks with more robustness and accuracy. In contrast, our work examines open models and tries to assess their fundamental capabilities. TableLlama Zhang et al. (2023) is similar to our work but focuses only on tabular data, which we consider a subset of SKG.
@@ 2.3 Reasoning Capability in LLMs
Reasoning stands as a pivotal skill for LLMs in the development of real-world AI applications which would enable the autonomous completion of many thought-intensive tasks viewed traditionally to require human thinking, like programming or mathematical problem-solving Li et al. (2022). Recent studies Li et al. (2022, 2023c); Rozière et al. (2023); Azerbayev et al. (2023a) indicate that LLMs trained on code and mathematical datasets exhibit profound reasoning skills, and can even achieve performance on par with human levels. For example, Code-Llama Rozière et al. (2023), a foundation model trained on more programming data, has significantly improved reasoning capabilities across a variety of programming and mathematical benchmarks. Furthermore, Llemma Azerbayev et al. (2023a) continues to pretrain the Code-Llama model on a mix of scientific papers, math-related web data, and mathematical code. Its results show excellent reasoning capabilities on the MATH benchmark Hendrycks et al. (2021) and the ability to prove theorems without further fine-tuning. On the fine-tuning side, WizardMath Luo et al. (2023a), and WizardCoder Luo et al. (2023c) have shown the effectiveness of instruction tuning on reasoning capabilities, given high quality data.
@@ 4 Experiments
Dataset Metric SoTA ChatGPT Base ST UL2 TabLLaMA USKG StructLM (Ours) $\Delta$
@@
KVRet All Micro 67.9 52.9 39.5 70.9 - 48.7 67.9 72.6 69.5 69.3 +4.5
HybridQA Acc 68.4 23.7 2.3 58.4 61.0 27.6 59.4 59.2 59.1 61.1 +1.7
@@
BIRD Acc 36.6 21.8 0.0 24.4 - 4.2 $\times$ 22.3 22.8 24.7 -
CoSQL EM 58.3 33.7 0.2 52.4 - 5.4 $\times$ 49.8 52.2 55.0 -
SQA Acc 70.5 18.7 2.3 60.4 70.1 2.0 $\times$ 49.7 36.1 44.2 -
Infotabs Acc 75.6 46.9 40.2 68.7 - 35.5 $\times$ 55.3 58.1 61.8 -
WikiTableText BLEU 33.7 3.8 5.7 39.8 - 10.2 $\times$ 8.3 9.3 8.8 -
Finqa Acc 71.1 31.4 1.7 79.7 - 18.6 $\times$ 27.3 25.6 36.2 -
Average 57.6 26.1 8.4 54.2 - 12.6 $\times$ 35.5 34.0 38.4 -
Table 2: The held-in evaluation results of our model against other baselines. TabLLaMA refers to the table fine-tuned LLaMA Zhang et al. (2023), the Held-out results are obtained by us. BASE refers to the 1-shot performance of CodeLlama-Instruct-7B. Single FT refers to the single-task finetuning result of CodeLlama-Instruct-7B on each dataset. XYZ denotes the state-of-the-art score achieved by us. All StructLM held-out results are 0-shot. Specifications as to how SoTA results are selected are given in Table 4.
@@
When compared to the single-task 18 USKG models, StructLM-7B can surpass USKG by a average of 2%. From a parameter-count perspective, each of the USKG models is a T5-3B model, which means over the entire held-in set, these results require 54B parameters. Our 7B StructLM in comparison can be viewed as being nearly 8x as parameter efficient while still surpassing USKG models on 11 of 18 datasets. It is worth noting that although our single-task (ST) models are more than double the size in parameters compared to USKG, they do not perform much better on average. This fact indicates that there may be significant unused model capacity that can be better utilized via more effective training regimes, such as our instruction tuning. Regarding FLAN-UL2-20B Tay et al. (2023b), which was also extensively trained on structure knowledge grounding tasks, StructLM outperforms it on 7 of the 9 mutually held-in datasets. Regarding TableLLaMA Zhang et al. (2023), which is a table-specialized LLM, StructLM can outperform it significantly across the board. Unlike StructLM, TableLLaMA cannot deal with structure beyond tables, which make it less generalizable than StructLM.
@@
On held out tasks, StructLM shows strong generalization performance, outperforming ChatGPT on 5 of 6 tasks. We note that our zero-shot prompting method for the FinQA Chen et al. (2021) dataset requires the model to generate custom code and is focused only on the financial domain, and the input to Infotabs Gupta et al. (2020a) is a different type of table structure unseen in the input data, with different labels for the boolean output categories than the choices from the held-in fact verification tasks. This generalization capability is non-existent in USKG models as the name of each held-in task is specified in their training inputs. When compared to TableLLaMA, StructLM also generalizes much better.
@@ Appendix A Appendix
GrailQA EM val 77.1 TIARA + GAIN (T5-3B) (Gu et al., 2023)
@@
Table 6: The evaluation results comparing finetuning done on different base models. Code refers to Codellama-Instruct-7B. Math refers to LLemma-7b. LLaMA refers to Llama-2-7b.

### 28 Feb 2024 v2

@@ Abstract
Structured data sources, such as tables, graphs, and databases, are ubiquitous knowledge sources. Despite the demonstrated capabilities of large language models (LLMs) on plain text, their proficiency in interpreting and utilizing structured data remains limited. Our investigation reveals a notable deficiency in LLMs’ ability to process structured data, e.g., ChatGPT lags behind state-of-the-art (SoTA) model by an average of 35%. To augment the Structured Knowledge Grounding (SKG) capabilities in LLMs, we have developed a comprehensive instruction tuning dataset comprising 1.1 million examples. Utilizing this dataset, we train a series of models, referred to as StructLM, based on the CodeLlama architecture, ranging from 7B to 34B parameters. Our StructLM series surpasses task-specific models on 14 out of 18 evaluated datasets and establishes new SoTA achievements on 7 SKG tasks. Furthermore, StructLM demonstrates exceptional generalization across 6 novel SKG tasks. Contrary to expectations, we observe that scaling model size offers marginal benefits, with StructLM-34B showing only slight improvements over StructLM-7B. This suggests that structured knowledge grounding is still a challenging task and requires more innovative design to push to a new level.
@@ 1 Introduction
The previous work Yu et al. (2020); Liu et al. (2021); Xie et al. (2022); zhang2023tablellama has been mostly focused on building task-specific models for different tasks with rather limited generalization ability. Building a generalist structure knowledge grounding (SKG) system across a wide range of tasks proves to be challenging. This is mainly due to the heterogeneity of data format and use cases. We evaluated GPT-3.5-Turbo Jiang et al. (2023) on 18 SKG tasks and observed that its performance is on average 35% lower than the SoTA specialized models. It shows that the LLM’s ability on SKG is heavily overlooked during the pre-training phase.
@@
In this paper, we explore the possibility of building a generalist model based on LLMs that can ground on diverse types of structure and unstructured knowledge to interface with humans. Specifically, we construct a large data set of over a million instruction-following examples, a majority of which is SKG data, along with additional general instruction-following data, which we find improves generalizability. We fine-tune models at three scales: 7B, 13B, and 34B, based on the CodeLlama family of code foundation models. When compared to USKG, we find that our 7B model surpasses these single-task models from on $11$ of $18$ tasks with our 34B model outperforming on $14$ of $18$. As depicted in Figure 1, StructLM achieves SoTA on $7$ out of 18 evaluated tasks, beating ChatGPT by a huge margin.
@@ 2.1 Solving SKG tasks
More recently, methods have begun to move away from these auxiliary task-specific structures. USKG Xie et al. (2022) were the first to unify many SKG tasks into a sequence-to-sequence format, allowing them to be aggregated into the same data mixture. StructGPT Jiang et al. (2023) represents a line of work that uses prompting frameworks on powerful LLMs to solve tasks with more robustness and accuracy. In contrast, our work examines open models and tries to assess their fundamental capabilities. Contemporary to our work, TableLlama Zhang et al. (2023) has argued that tabular data deserves special attention. Focusing on this domain, their method fine-tunes on several new tabular tasks to improve table understanding, and operates on a longer 8k context length. These improvements can be additive to our work.
@@ 2.3 Reasoning Capability in LLMs
Reasoning stands as a pivotal skill for LLMs in the development of real-world AI applications which would enable the autonomous completion of many thought-intensive tasks viewed traditionally to require human thinking, like programming or mathematical problem-solving Li et al. (2022). Recent studies Li et al. (2022, 2023c); Rozière et al. (2023); Azerbayev et al. (2023a) indicate that LLMs trained on code and mathematical datasets exhibit profound reasoning skills, and can even achieve performance on par with human levels. For example, CodeLlama Rozière et al. (2023), a foundation model trained on more programming data, has significantly improved reasoning capabilities across a variety of programming and mathematical benchmarks. Furthermore, Llemma Azerbayev et al. (2023a) continues to pretrain the CodeLlama model on a mix of scientific papers, math-related web data, and mathematical code. Its results show excellent reasoning capabilities on the MATH benchmark Hendrycks et al. (2021) and the ability to prove theorems without further fine-tuning. On the fine-tuning side, WizardMath Luo et al. (2023a), and WizardCoder Luo et al. (2023c) have shown the effectiveness of instruction tuning on reasoning capabilities, given high quality data.
@@ 4 Experiments
Dataset Metric SoTA ChatGPT Base ST FLAN-UL2 TableLlama USKG StructLM (Ours) $\Delta$
@@
KVRet All Micro 67.9 52.9 39.5 70.9 - - 67.9 72.6 69.5 69.3 +4.5
HybridQA Acc 68.4 23.7 2.3 58.4 61.0 - 59.4 59.2 59.1 61.1 +1.7
@@
BIRD Acc 36.6* 21.8 0.0 24.4* - - - 22.3 22.8 24.7 -
CoSQL EM 58.3* 33.7 0.2 52.4* - - - 49.8 52.2 55.0 -
SQA Acc 70.5* 18.7 2.3 60.4* 70.1* - - 49.7 36.1 44.2 -
Infotabs Acc 75.6* 46.9 40.2 68.7* - 35.5 - 55.3 58.1 61.8 -
WikiTableText BLEU 33.7* 3.8 5.7 39.8* - 10.2 - 8.3 9.3 8.8 -
Finqa Acc 71.1* 31.4 1.7 79.7* - 18.6 - 27.3 25.6 36.2 -
Average 57.6* 26.1 8.4 54.2* - - - 35.5 34.0 38.4 -
Table 2: The overall evaluation results of our model against other baselines. Cells with "-" in the held-in part mean that the model was trained on these datasets, thus the results are not comparable. Cells with "-" in the held-out part mean the results are not yet available. Cells in the held-out section with "*" are held-in results and other than ST, are copied from the original papers for reference. TableLlama held-out results are obtained by running their checkpoint. BASE refers to the 1-shot performance of CodeLlama-Instruct-7B. ST refers to the single-task finetuning result of CodeLlama-Instruct-7B on each dataset. XYZ denotes the state-of-the-art score achieved by us. All StructLM held-out results are 0-shot. Specifications as to how SoTA results are selected are given in Table 4.
@@
When compared to the single-task 18 USKG models, StructLM-7B can surpass USKG by a average of 2%. From a parameter-count perspective, each of the USKG models is a T5-3B model, which means over the entire held-in set, these results require 54B parameters. Our 7B StructLM in comparison can be viewed as being nearly 8x as parameter efficient while still surpassing USKG models on 11 of 18 datasets. It is worth noting that although the single-task (ST) models are more than double the size in parameters compared to USKG, they do not perform much better on average. This fact indicates that there may be significant unused model capacity that can be better utilized via more effective training regimes, such as our instruction tuning. Regarding FLAN-UL2-20B Tay et al. (2023b), which was also extensively trained on structure knowledge grounding tasks, StructLM outperforms it on 7 of the 9 mutually held-in datasets. Our results on held-in datasets (Tabfact and FeTaQA) are on par with TableLlama Zhang et al. (2023), which is an LLM pre-trained on 2.6M table understanding tasks.
@@
On held out tasks, StructLM shows strong generalization performance, outperforming ChatGPT on 5 of 6 tasks. We note that our zero-shot prompting method for the FinQA Chen et al. (2021) dataset requires the model to generate custom code and is focused only on the financial domain, and the input to Infotabs Gupta et al. (2020a) is a different type of table structure unseen in the input data, with different labels for the boolean output categories than the choices from the held-in fact verification tasks. This generalization capability is non-existent in USKG models as the name of each held-in task is specified in their training inputs.
@@ Appendix A Appendix
GrailQA EM val 77.1 TIARA + GAIN (T5-3B) (Shu and Yu, 2023)
@@
Table 6: The evaluation results comparing finetuning done on different base models. Code refers to CodeLlama-Instruct-7B. Math refers to Llemma-7b. LLaMA refers to Llama2-7b.